OpenStudy (anonymous):

I need stats help, y'all. For a sample with a normal distribution, n=12 and s=4.2. The null hypothesis is σ≤3.0. Could someone help me get started?

OpenStudy (inkyvoyd):

I have no idea what you're talking about, but NULL HYPOTHESIS IS A LOGICAL FALLACY. ALso, @FoolForMath

OpenStudy (zarkon):

you need to reject \(H_0\) if \(\displaystyle\frac{(n-1)s^2}{\sigma_0^2}\le \chi^2_{1-\alpha}\) with n-1 d.f.